# Scalable, Tokenization-Free Diffusion Model Architectures with Efficient Initial Convolution and Fixed-Size Reusable Structures for On-Device Image Generation

### v1

@@ 2.1 Diffusion Models
[…]
@@ 3.2 Implementation details
[…]
@@
Core structure configaration: We conducted experiments by concatenating $N$ core structures, with $N={12,24,32}$ core blocks. Increasing the number of such blocks (making the model deeper) and increasing the embedding dimension $L$ (making the model wider) improves model performance in terms of the FID score, as shown in Figure 3(d). However, this improvement comes with an increase in the complexity, as shown in Figure 3(a).
Figure 5: Text conditional image generation on MS-COCO. Text prompts: 1. A green train is coming down the tracks. 2. A group of skiers are preparing to ski down a mountain. 3. A small kitchen with a low ceiling. 4. A group of elephants walking in muddy water. 5. A living area with a television and a table. 6. A road with traffic lights, street lights and cars. 7. A bus driving in a city area with traffic signs. 8. A bus pulls over to the curb close to an intersection. 9. A group of people are walking and one is holding an umbrella. 10. A baseball player taking a swing at an incoming ball.
@@ 4.2 Main result
Text conditional image synthesis: We evaluated the performance of our STOIC-based diffusion model for text-conditional image synthesis on the MSCOCO dataset, comparing it against various benchmarks, as shown in Table 2. The text description of the MSCOCO dataset is encoded using the text encoder of CLIP [31]. This embedding is subsequently fed to our model for text-conditional image synthesis. We employed a pretrained autoencoder from LDM [34] to encode images of size $256\times 256\times 3$ into latent representations of size $32\times 32\times 4$ for the MSCOCO dataset. We conducted experiments with two variants, STOIC-$\mathcal{S}$-1 and STOIC-$\mathcal{S}$-2, and with different numbers of model parameters. Notably, STOIC-$\mathcal{S}$-2 achieves an FID of 8.69 with 101M parameters, outperforming GAN-based models such as VQ-Diffusion (370M parameters) and XMC-GAN (166M parameters), despite having fewer parameters. In Figure 5, we present the results of our text-to-image synthesis model across various text prompts. We observed that the STOIC-based diffusion model, suitable for deployment on mobile devices for inference, generates higher-quality samples with improved alignment between the text and the generated images. In contrast, the UViT model requires higher GMAC (with our GMAC count being 22.35 compared to UViT’s 28) for comparable performance and incurs high latency on hardware devices due to long skip connections.

### v2

# Hardware-Friendly Diffusion Models with Fixed-Size Reusable Structures for On-Device Image Generation

@@ 2.1 Diffusion Models
$\underset{\theta}{\arg\min}\mathbb{E}\lvert\lvert\epsilon-\epsilon_{\theta}(x_{t},t)\rvert\rvert_{2}^{2}$ (1)
@@ 3.2 Implementation details
$H_{o}=\lfloor\frac{H_{i}+2P-\mathcal{K}-1}{\mathcal{S}}+1\rfloor\>,\>W_{o}=\lfloor\frac{W_{i}+2P-\mathcal{K}-1}{\mathcal{S}}+1\rfloor$, where $\lfloor\cdot\rfloor$ denotes the floor function, which returns the greatest integer less than or equal to its argument. Hence, with $\{\mathcal{S}=1,\mathcal{K}=3,P=1\}$, the output dimension $H_{o}=H_{i}$ and $W_{o}=W_{i}$. With $\{\mathcal{S}=2,\mathcal{K}=2,P=0\}$, the output dimension $H_{o}=\frac{H_{i}}{2}$ and $W_{o}=\frac{W_{i}}{2}$.
@@
Core structure configaration: We conducted experiments by concatenating $N$ core structures, with N=12,24,32N={12,24,32} core blocks. Increasing the number of such blocks (making the model deeper) and increasing the embedding dimension $L$ (making the model wider) improves model performance in terms of the FID score, as shown in Figure 3(d). However, this improvement comes with an increase in the complexity, as shown in Figure 3(a).
Figure 5: Qualitative comparison between (a) STOIC and (b) U-ViT for text-conditional image generation on the MS-COCO dataset. STOIC-$\mathcal{S}-2$ maintains a comparable parameter count ($\sim$43M) to U-ViT, while achieving significantly lower computational complexity, with approximately 11 GMACs compared to U-ViT’s 43 GMACs. Text prompts: 1. A train is coming down the tracks. 2. A group of skiers are preparing to ski down a mountain. 3. A small kitchen with a low ceiling. 4. A group of elephants walking in muddy water. 5. A living area with a television and a table. 6. A road with traffic lights, street lights and cars. 7. A bus driving in a city area with traffic signs. 8. A bus pulls over to the curb close to an intersection. 9. A group of people are walking and one is holding an umbrella. 10. A baseball player taking a swing at an incoming ball.
@@ 4.2 Main result
Text conditional image synthesis: We evaluated the performance of our STOIC-based diffusion model for text-conditional image synthesis on the MSCOCO dataset, comparing it against various benchmarks, as shown in Table 2. The text description of the MSCOCO dataset is encoded using the text encoder of CLIP [31]. This embedding is subsequently fed to our model for text-conditional image synthesis. We employed a pretrained autoencoder from LDM [34] to encode images of size $256\times 256\times 3$ into latent representations of size $32\times 32\times 4$ for the MSCOCO dataset. We conducted experiments with two variants, STOIC-$\mathcal{S}$-1 and STOIC-$\mathcal{S}$-2, and with different numbers of model parameters. Notably, STOIC-$\mathcal{S}$-2 achieves an FID of 8.69 with 101M parameters, outperforming GAN-based models such as VQ-Diffusion (370M parameters) and XMC-GAN (166M parameters), despite having fewer parameters. In Figure 5, we present the results of our text-to-image synthesis model across various text prompts with UViT [1]. We observed that the STOIC-based diffusion model, suitable for deployment on mobile devices for inference, generates higher-quality samples with improved alignment between the text and the generated images. In contrast, the UViT model requires higher GMAC (with our GMAC count being 22.35 compared to UViT’s 28) for comparable performance and incurs high latency on hardware devices due to long skip connections.